OpenStudy (anonymous):

What is the solution to the equation 5x + 2(x − 4) = 5x + x − 10? 3 2 −2 −3

OpenStudy (anonymous):

The steps below show the incomplete solution to find the value of x for the equation 5x − 3x + 4 = −4 + 20: Step 1: 5x − 3x + 4 = −4 + 20 Step 2: 5x − 3x + 4 = 16 Step 3: 2x + 4 = 16 Which of these is most likely the next step? 2x = 4 2x = 12 2x = 20 2x = 18